My son has this math problem and I don't remember for the life of me how to do this. Could you show how to get the answer, so I can help him with others.
3x - 4y = 5
6x - 8y = -5

Answer by edjones(8007)

3x - 4y = 5
6x - 8y = -5
.
If you multiply the top equation by 2 you get:
6x-8y=10
Since 6x-8y can't equal 10 and -5 then the values of x and y are undefined (there is no answer).
.
Ed